Write a rational function satisfying the following criteria:
Vertical assymptote: x=1
Assymptote: y=x+1
zero of the function: x=-2

Answer by stanbon(75887)

Write a rational function satisfying the following criteria:
Vertical asymptote: x=1
so f(x) must have a factor of (x-1) in the denominator
-------------------
Asymptote: y=x+1
So, f(x) = g(x)/(x-1) = x+1 + k/(x-1)
Need to find g(x) and k
--------
zero of the function: x=-2
so f(x) must have a factor of (x+2) in the numerator.
--------------------------------
Working the problem:
f(x) = [(x+2)(ax+b)/(x-1)] = x+1 + k/(x-1)
Then:
f(-2) = 0 = -2+1+k/(-2-1)
k/-3 = 1
k = -3
---------------
f(x) = [(ax^2+(b+2a)x+2b)/(x-1)] = x+1 -3/(x-1)
Need to find a and b:
Divide ax^2+(b+2a)x+2b by (x-1) to force a quotient of (x+1) and remainder =-3:
This will happen if a=1 and b = -2
----------------
So, f(x)= [(x+2)(x-2)/(x-1)]
==================
Checking:
(x^2-4)/(x-1) = x+1 -3/(x-1) which has a vertical
asymptote at x=1, a zero at x=-2, and an asymptote of x+1
===================
Cheers,
Stan H.